OpenStudy (anonymous):

This is super hard.. Can anyone help me please?Before a football game, a coin toss is used to determine which team will kick off. The height h (in feet) of a coin above the ground t seconds after being flipped up into the air is given by h= -16t^2 + 12t +4 . How long does a team captain have to call heads or tails if it must be done while the coin is in the air?

OpenStudy (anonymous):

it's not that hard... without all those words, you'll end up having to solve this equation: \[\large 0=-16t^2+12t+4 \] can you solve that equation for t?

OpenStudy (anonymous):

@dpaInc 4(4t^2 + 3t + 1) ?

OpenStudy (anonymous):

you factored a 4 or -4?

OpenStudy (anonymous):

whoops -4

OpenStudy (anonymous):

let's factor out the -4

OpenStudy (anonymous):

\[\large -4(4t^2-3t-1)=0 \] \[\large 4t^2-3t-1=0 \]

OpenStudy (anonymous):

Ok I got that so far

OpenStudy (anonymous):

now you can either factor this (which is possible, btw) or if you don't want to think, just use the quadratic formula....

OpenStudy (anonymous):

what do you wanna do?

OpenStudy (anonymous):

I only know how to factor how would I use the formula though?

OpenStudy (anonymous):

t=-1/4 t=1 ?

OpenStudy (anonymous):

ok... before using the formula, the equation needs to be in standard from so you can identify the coefficients a, b, and c: standard form: \(\large at^2+bt+c=0 \) quadratic formula: \(\large t=\frac{-b\pm \sqrt{b^2-4ac}}{2a} \)

OpenStudy (anonymous):

@dpaInc I got t= -1/4 t=1 So it's wanting to know... How long does a team captain have to call heads or tails if it must be done while the coin is in the air? And the answers are in seconds?...

OpenStudy (anonymous):

in your equation, a=4 b=-3 c=-1 plug these values into the formula for your answer... notice you'll get two values... choose the one that makes sense in the real world... choose the positive t...

OpenStudy (anonymous):

yes... seconds....

OpenStudy (anonymous):

so @dpaInc 1 second?

OpenStudy (anonymous):

@dpaInc Is 1 sec correct?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

Why wouldn't it be 0.25 sec? @dpalnc

jimthompson5910 (jim_thompson5910):

Your other solution was t = -1/4, but that doesn't make sense in this context because you can't have a negative time.

jimthompson5910 (jim_thompson5910):

So you ignore that solution

OpenStudy (anonymous):

Oh ok I get it thanks so much!!!!! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome